FULLY STATE-RESOLVED PHOTODISSOCIATION OF FORMALDEHYDE: H_{2}CO → H + HCO. K-CONSERVATION AND A RIGOROUS TEST OF STATISTICAL THEORIES

HONGMING YIN, KLAAS NAUTA and \underline{SCOTT H. KABLE}, School of Chemistry, University of Sydney, NSW, 2006, Australia.

The photodissociation dynamics of the reaction H_{2}CO + hν → H + HCO have been investigated in the range 60-400 cm^{−1} above the reaction threshold. Supersonically-cooled formaldehyde was excited into 15 specific J, K_{a}, K_{c} rotational states in two vibrational levels 2^{1}4^{1}6^{1} and 2^{2}4^{1} in the Ã(^{1}A_{2}) state. The laser induced fluorescence spectra of the nascent HCO fragment provided detailed product state distributions (PSDs), resolved by N, K_{a}, K_{c} and J. When just the overall molecular rotation, N, is considered the PSDs are in remarkable agreement with calculations based on phase space theory (PST). However, when the projection of N onto the molecular frame (K_{a}, K_{c}) is included the distributions show consistent deviations from PST. In particular, there is a tendency to preserve the initial parent rotational motion about the a and b axes. The effect is that states with higher initial K_{a} in H_{2}CO produce higher final K_{a} in the HCO fragment. There is also a tendency for the upper/lower members of the asymmetry doublets in H_{2}CO to map onto the same upper/lower set of product state asymmetry doublets. Finally, there are oscillations in some of the detailed PSDs that remain unexplained.